Determine the domain of each of the following functions: 1. f(x) = x + 10 6. A(x) = x2 -2 2. F(x) = 2 3 π‘₯ + 5 7. H(x) = √π‘₯ βˆ’ 2 3. g(x) = 5 – 3x 8. K(x) = √π‘₯ 2 βˆ’ 2 4. g(x) = 1 (π‘₯+5)(π‘₯βˆ’1) 9. C(x) = 2x3 + 4x2 - 2x + 1 5. b(x) = π‘₯βˆ’1 π‘₯ 2+5π‘₯+6 10. √π‘₯+1 π‘₯βˆ’2Β 


1
Expert's answer
2021-12-09T16:38:20-0500

The domain of the function(D) is the set of all values that argument might take. I will assume that x is a real number. Also the conditions is inaccurate, so i don't fully sure whether i recognized it correctly in each case, but it must be close to it

f(x) = x + 10. "D:x\isin R"


"A(x) = x^2 -2" . "D:x\isin R"


"F(x) = 2^{3\ud835\udc65} + 5" . "D:x\isin R"


"H(x) = \sqrt𝑥 − 2" . The value under the square root must be non-negative, so "D:x\isin [0,+\infty)"


g(x) = 5 – 3x . "D:x\isin R"


"K(x) = \sqrt{x^2 − 2}" . "D:x^2-2≥0\implies x^2≥2\implies D:x\isin (-\infty,-\sqrt2)\cup(\sqrt2,+\infty)"


"g(x) = {\frac 1 {(𝑥+5)(𝑥−1)}}" . Cannot divide by 0, so "D:(x+5)(x-1)\not=0\implies D:x\in R" \ {-5, 1}


"C(x) = 2x^3 + 4x^2 - 2x + 1" . "D:x\isin R"


"b(x) = 𝑥−{\frac 1 {x^2+5x+6}}" . "D:x^2+5x+6\not=0\implies D:x\in R" \ {-3, -2}


"f(x)={\frac {\sqrt{x-1}} {x-2}}" . "D:(x-1≥0)\land (x-2\not=0)\implies D:x\isin [1,+\infty)" \ {2}